OpenStudy (anonymous):

What's the missing no.? 1, 2, 7, ____ 157, 782

OpenStudy (cwrw238):

up 1 = 2 up 1*5 = 7 up 5*5 = 32 up 5*5*5 = 157 up 5^4 = 782 missing number is 32

OpenStudy (anonymous):

Thanks!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

what do you mean by "up"?

OpenStudy (cwrw238):

add